# Jie Wu1∗  Haoling Li1∗  Xin Zhang2∗†‡  Jiani Guo3  Jane Luo2  Steven Liu2 Yangyu Huang2Ruihang Chu1‡Scarlett Li2Yujiu Yang1

### v1

@@ 5 Discussion
Test-time Scaling. We compare the pass@k performance of Qwen2.5-Coder-7B-Instruct, Qwen3-8B, X-Coder-7B-SFT, and X-Coder-7B in Figure 8. X-Coder-7B outperforms its foundation model by 51.3 points in pass@16, and matches Qwen3-8B with 8× fewer rollouts. Moreover, X-Coder shows a larger gap between pass@1 and pass@16 compared to Qwen3-8B (19.2 vs. 13.8), indicating greater diversity in the reasoning patterns it can explore. Although RL models begin with higher initial performance than the SFT model, the gap does not expand within 16 rollouts, suggesting that RL improves pass@1 but may not escape its starting point (Wu et al., 2025).
Behaviors after SFT and RL. After SFT, the model frequently exhibits cognitive behaviors such as planning, verification, backtracking, and reflection, as illustrated by the case study in §H.1. This suggests that such behaviors can be directly distilled from the teacher rather than induced by the RL process. During the later stages of RL, the model shows signs of reward hacking, attempting to exploit edge cases for partial rewards instead of producing genuine solutions, as detailed in §H.3. We also observe several bad patterns in code reasoning, including premature termination when the model is aware that the context is running out, recalling memorized submissions in C++ and attempting to translate them into Python, and emitting incomplete code before the context window is exhausted. These cases are illustrated in §H.2.
6 Related Work
Data Synthesis for Code. The research community has long recognized the scarcity of high-quality coding tasks. To address this, Wizard-Coder (Luo et al., 2024) extends Evol-Instruct (Xu et al., 2024) by evolving basic code-instruction data into augmented variants. rStar-Coder (Liu et al., 2025a) further adapts this augmentation strategy to the competitive programming domain. CodeEvo (Sun et al., 2025) introduces a coder–reviewer interaction framework to collaboratively synthesize high-quality instruction–code pairs.
SelfCodeAlign (Wei et al., 2024) advanced task synthesis beyond simple seed evolution by introducing concept composition. It extracts fundamental concepts from seed problems and uses these building blocks to generate a vastly larger bank of novel coding tasks through the systematic combination of underlying concepts, pioneering a new pathway for scaling problem synthesis. Epicoder (Wang et al., 2025) follows this direction by sampling sub-features from a large and expressive feature tree to formulate novel problems, further improving task complexity and diversity.
This work targets competitive programming, a domain where previous synthetic methods struggle to synthesize tasks requiring deep reasoning and accurate test cases. We demonstrate that a fully synthetic pipeline offers a practical and scalable solution to these challenges. Methodologically, we employ competition-oriented feature extraction to synthesize challenging, coherent, and diverse tasks. Crucially, we construct high-fidelity test cases using prompt- and tool-based input generation combined with voting-based labeling. These accurate tests enable Code RL advancements.
Post-training Recipe for Code Reasoning Model. From the training perspective, current approaches to building coding-expert LLMs generally fall into three paradigms: (i) purely supervised fine-tuning on real-world tasks or their rewritten or evolved variants (Labs, 2025; Guha et al., 2025; Liu et al., 2025a), (ii) purely reinforcement-based fine-tuning using a GRPO-related (Shao et al., 2024b; He et al., 2025; Luo et al., 2025; Fu et al., 2025) algorithm, and (iii) reinforcement learning staged after supervised fine-tuning on mixed coding and mathematical data (Liu et al., 2025b; Xiaomi et al., 2025; Su et al., 2025). High-quality code data is scarcer than mathematical data. Consequently, existing approaches rely heavily on real-world data and lack a stable two-stage recipe for coding expertise, often mixing in mathematics with little evidence of success on code alone. In this paper, we show that stable and consistent improvements in code reasoning can be achieved solely with synthetic data, while also reducing the risk of data leakage shown in §G.
@@ 7 Conclusion
In this paper, we explore a fully synthetic approach to competitive programming and propose a novel data synthesis framework that demonstrates how synthetic tasks, solutions, and tests can train large reasoning models to achieve significant performance gains, thereby reducing reliance on real-world data. Building on this framework, we contribute scalable synthetic SFT and RL training sets, supported by a dedicated RL infrastructure, and introduce the X-Coder series. Furthermore, we provide insights into code-centric SFT-then-RL training, ablate key factors that shape performance, and present in-depth analyses with illustrative case studies of code reasoning models.
@@ Appendix B Novel Task Synthesis
First, rather than relying on broad feature definitions, we explicitly extract and evolve competition-related features from 10,000 question–solution pairs in TACO (Li et al., 2023) using GPT-4o-0513 (§B.1). Second, we adopt a two-stage process: selecting mutually consistent features and then formulating challenging, hint-free tasks (§B.2). Third, we extend the synthesis method to support multi-style generation, covering CodeForces-style tasks (rich narratives with standard I/O), LeetCode-style tasks (starter code with fixed signatures), and AtCoder-style tasks (concise specifications), thereby enhancing task diversity. In §B.3, we further estimate the difficulty of synthesized problems using a trained discriminator.
@@ B.1 Feature Extraction and Evolution
While EpiCoder extracts general-purpose features from raw corpus, we explictly extract and evol compatetitive programming-related feature. Speficilly, we design multiple aspect of features that highly relates to competitve programming, such as data structure, algorithm, mathmatical, ect.
@@
"data structures": [
"array",
"linked list",
"heap",
"segment tree"
@@
3.3 Dynamic Programming: Specify dynamic programming techniques. Example: ["Longest Increasing Subsequence", "Knapsack Problem"].
3.4 Search Algorithms: Identify search algorithms used. Example: ["Binary Search", "Linear Search"].
3.5 Other relevant subcategories...
@@
8. Purpose: What the code is used to do. Example: "To find the shortest path in a graph using Dijkstras algorithm."
9. Summary: Provide a concise summary. Example: "Solves the given competitive programming problem using a depth-first search approach to traverse the graph."
@@
You are provided with a feature tree represented as a nested JSON structure. Each node in this tree represents a feature or a sub-feature of competitive algorithm programming, with the leaves being the most specific features. Your task is to expand this feature tree both in depth and breadth. Depth expansion means adding more specific sub-features to existing leaves. Breadth expnasion means adding more sibling features at the current levels.
@@
"mathematics": {
"number theory": [
"prime factorization",
"greatest common divisor",
"power modular reduction"
"combinatorics": [
"Pascals triangle",
"permutations and combinations",
"binomial coefficients"
@@ B.1.1 Statistics for Feature Extraction and Evolution
We present detailed statistics on feature evolution and data filtering to demonstrate how the pipeline expands feature diversity and yields a high-quality 240k dataset. The statistics of feature extracted and evoled as follows.
@@
The evolution strategy greatly enhances both the quantity and diversity of features, providing support for generating diverse tasks.
@@ B.2 Stylized Task Generation for Competitive Programming
Stage 1 Prompt Template for Feature Selection
STAGE1_PROMPT_TEMPLATE = """You are a professional competitive programming problem setter.
Your task consists of three parts:
@@
"iterative binary search": {{
"potential_use": "Efficient loop-based implementation for finding bounds or specific elements."
"breadth-first search (BFS)": {{
"level-order BFS": {{
"potential_use": "Traverses graphs layer by layer; useful for shortest path or component discovery."
"data structures": {{
"bitmap": {{
"bit manipulation": {{
"bitwise AND": {{
"potential_use": "Filters or checks properties using bitmasks."
"bitwise OR": {{
"potential_use": "Combines flags or sets with bitwise aggregation."
@@
"data structures": {{
"bitmap": {{
"bit manipulation": {{
"bitwise AND": {{
"feature": "bitwise AND",
"potential_use": "Filters or checks properties using bitmasks."
@@
Instructions:
- Always preserve the tree structure in "feature_roles_tree" and "selected_features_tree".
- In selected_features_tree, only include "feature" and "potential_use" fields for leaf nodes.
- "integration_strategy" should make clear how/why these features form a coherent, advanced problem.
- Do not be overly conservative; it is often possible to design advanced problems where many features interact in non-trivial ways. Challenge yourself to maximize feature use without sacrificing problem quality.
@@ B.2.1 Compatibale Feature Selection
We present a case to examine how model selects compatibale features and combine them.
@@
"string_algorithms": {
"pattern_matching": [
"KMP algorithm",
"Boyer-Moore"
"data_structures": {
"tree_structures": [
"segment tree",
"fenwick tree"
"hash_structures": [
"rolling hash",
"cuckoo hashing"
"optimization_techniques": {
"dynamic_programming": [
"interval DP",
"tree DP"
LLM pairs each feature with potentially usage to obtain feature tree with role annotation. For example, LLM will anonotes feature “rolling hash” as “Compute hash values for sliding windows in constant time”. These annotations help LLM to aggregate these features based on their potentially usage. For above given feature tree, the feature tree with potential usage looks like:
@@
"network_flow": {
"Ford_Fulkerson": {
"potential_use": "Find maximum flow using augmenting paths with DFS"
"Edmonds_Karp": {
"potential_use": "Guarantee polynomial time maximum flow using BFS"
"string_algorithms": {
"pattern_matching": {
"KMP_algorithm": {
"potential_use": "Linear time pattern matching with failure function preprocessing"
"Boyer_Moore": {
"potential_use": "Efficient pattern search using bad character and good suffix heuristics"
"data_structures": {
"tree_structures": {
"segment_tree": {
"potential_use": "Support range queries and updates in O(log n) time"
"fenwick_tree": {
"potential_use": "Efficiently compute prefix sums with point updates"
"hash_structures": {
"rolling_hash": {
"potential_use": "Compute hash values for sliding windows in constant time"
"cuckoo_hashing": {
"potential_use": "Achieve worst-case O(1) lookup with multiple hash functions"
"optimization_techniques": {
"dynamic_programming": {
"interval_DP": {
"potential_use": "Solve optimization problems on contiguous ranges"
"tree_DP": {
"potential_use": "Compute optimal solutions on tree structures bottom-up"
LLM then selects a compatible and consistent subtree that can formulate a self-contained compatetiting programming problem. For example, LLM keeps the features that can be aggregated into selected feature tree, and concluding with an intergration strategy, which displays how to combine these features into a unified problem.
@@
"optimization_techniques": {
"dynamic_programming": {
"tree_DP": {
"feature": "tree DP",
"potential_use": "Optimize subproblems on network tree decomposition"
@@ B.2.2 From Feature to Stylized Task
During task generation, LLM recieves selected features tree and its integration strategy to formulate styleized task based on prompt recieved. Task generation prompt for Codeforces-style is as follows:
"""You are a professional competitive programming problem setter.
You have been provided with:
- selected_features_tree: a tree structure where each leaf contains a "feature" name and its "potential_use".
- integration_strategy: a strategy describing how these features should be integrated into a single, high-quality problem.
Your task is to **generate a complete Codeforces-style problem statement** that fully integrates ALL selected features.
Requirements:
- The story and setting must naturally motivate every selected feature, making each indispensable for an optimal solution.
- Specify precise input/output format and tight constraints.
- Provide at least two distinct, non-trivial sample Input/Output pairs, each with a clear explanation.
- Make sure the samples are consistent with your constraints and the solution requires use of all selected features.
- Do not include any references to algorithms, data structures, solution strategies, or any implicit or explicit hints in any part of the statement, notes, or examples. Do not include any motivational, summary, or instructional phrases (e.g., "Remember", etc.) at any point in the output. The statement must end after the final example or clarification, with no extraneous commentary.
- Output should be a **single JSON object** with the field "question" only.
**Output Format (strictly):**
[…]
**Inputs:**
- selected_features_tree (JSON):
{selected_features_info}
- integration_strategy (string):
{integration_strategy}
Instructions:
- You must ensure every selected feature is essential and naturally integrated.
- Output ONLY the required JSON object, no extra text.
In this instance, our generated Codeforces problem is shown in Figure 14, while the generated AtCoder and LeetCode problems are presented in Figures 15 and 16, respectively.
@@
Figure 15: Case for AtCoder-style Problem, featuring concise, minimal explainations.
Figure 16: Case for LeetCode-style Problem, featuring predefined function signatures.
@@
The rationale for above two-stage pipeline is that a single-step approach is less effective. When performing both steps simultaneously, LLMs tend to oversimplify complex instructions into trivial cases, reducing both diversity and difficulty of the generated task.
To empirically validate this, we generated 32k tasks using the one-step method (feature-tree $\rightarrow$ task) and using proposed “two-stage” method (feature-tree $\rightarrow$ sub-tree $\rightarrow$ task). The SFT results on LiveCodeBench v5 are as Table 9. The 5.3 gain shows that explicit sub-tree selection and integration is significantly helpful for producing high-quality, challenging tasks and justifies SynthSmith’s modular design.
@@ B.3 Task Difficulty Estimates
Judging the difficulty of a synthetic task is challenging. To better capture the difficulty distribution of tasks generated by X-Coder, we adopt a classifier-based approach. Specifically, we add a special classification token to Qwen2.5-Coder-14B-Instruct and fine-tune it to predict the Codeforces rating of 6,246 tasks from the CodeContests dataset with annotated ratings, reserving 5% as a validation set. The fine-tuned model achieves 84% classification accuracy on the validation set. We then use this model to estimate the difficulty of 1,000 tasks generated by our pipeline, obtaining a holistic distribution as shown in Table 10.
Table 10: Difficulty distribution of Codeforces-style ratings. “Original” denotes the annotated distribution from CodeContests, and “Test” denotes 1,000 tasks generated by our pipeline.
@@ D.1 Prompting-based Test Generation
You are a professional test case generation expert, skilled at designing comprehensive test cases for programming problems.
Please generate 15 different test cases for the following programming problem, including edge cases, small-scale, medium-scale, and large-scale test data.
@@ 5. Ensure test cases can thoroughly validate the correctness of solutions
D.2 Tool-based Test Generation
@@ 8. The program should use argparse to provide configurable random seed control:
Input/Output (IO)
The IO library helps you easily create test data files.
Constructor Options:
‘‘‘python
# Basic file specification
IO("test1.in", "test1.out") # Explicit input/output files
IO(file_prefix="test") # Generates test.in and test.out
IO(file_prefix="test", data_id=3) # Generates test3.in and test3.out
# Advanced file naming
IO(file_prefix="test", data_id=6,
input_suffix=".input", output_suffix=".answer") # test6.input and test6.answer
# Partial output options
IO("test2.in") # Only input file, output goes to temporary file
IO(file_prefix="test", data_id=5, disable_output=True) # No output file generated
IO() # Both files temporary (for use with comparator)
Note: Combine file_prefix and data_id with loops for batch generation.
IO Methods:
‘‘‘python
io = IO("test1.in", "test1.out") # Initialize IO object
# Input writing methods
io.input_write(1, 2, 3) # Writes "1 2 3" to input file (no newline)
[…]
io.input_write([1, 2, 3]) # Writes list as space-separated "1 2 3"
io.input_write(1, 2, 3, separator=’,’) # Writes "1,2,3," (note: current version leaves trailing comma)
# Output writing methods
io.output_write(1, 2, 3) # Writes "1 2 3" to output file
[…]
io.output_write(1, 2, [1, 2, 3], [4]) # Flattens nested lists to "1 2 1 2 3 4"
# Program execution
io.output_gen("~/Documents/std") # Runs program with input, captures stdout as output
[…]
Graph Generation
The Graph library generates various graph structures.
Manual Construction:
‘‘‘python
# Graph initialization
graph = Graph(10) # 10-node undirected graph (nodes 1-10)
graph = Graph(10, directed=True) # Directed version
# Adding edges
graph.add_edge(1, 5) # Default weight=1
graph.add_edge(1, 6, weight=3) # Custom weight
# Edge access and properties
graph.edges # Adjacency list containing Edge objects
for edge in graph.iterate_edges():
edge.start # Source node
edge.end # Target node
edge.weight # Edge weight
# Output formatting options
io.input_writeln(graph) # Default "u v w" per line
io.input_writeln(graph.to_str(shuffle=True)) # Random edge order
io.input_writeln(graph.to_str(output=Edge.unweighted_edge)) # "u v" format
Template Graphs:
‘‘‘python
# Basic graph templates
Graph.graph(n, m) # n nodes, m edges (weight=1)
Graph.graph(n, m, directed=True, weight_limit=(5, 300)) # Directed with weight range
Graph.graph(n, m, self_loop=False, repeated_edges=False) # No duplicate edges
# Special graph types
Graph.chain(n) # n-node chain (alias for tree(n, 1, 0))
Graph.flower(n) # n-node star graph (alias for tree(n, 0, 1))
Graph.tree(n) # Random tree
Graph.tree(n, 0.4, 0.35) # 40% chain-like, 35% star-like, 25% random
Graph.binary_tree(n) # Random binary tree
# Competition-specific graphs
Graph.hack_spfa(n) # Graph that breaks SPFA (1.5n edges)
Graph.hack_spfa(n, extra_edge=m) # With additional edges
Graph.DAG(n, m) # Directed Acyclic Graph
Graph.UDAG(n, m) # Undirected Connected Graph
Note: Most templates support weight_limit‘, weight_gen‘, self_loop‘, and repeated_edges parameters.
Polygon
Generate and analyze polygons.
‘‘‘python
# Polygon creation (points must be ordered)
p = Polygon([(0,0), (0,4), (4,4), (4,0)]) # Rectangle
# Geometric properties
p.perimeter() # Calculates perimeter
p.area() # Calculates area
# Generation templates
Polygon.convex_hull(n) # n-point convex hull
Polygon.simple_polygon(n) # Simple polygon (non-intersecting)
Vector
Generate unique vectors/number sequences.
‘‘‘python
# Basic usage
Vector.random() # Default: 5 unique numbers in [0,10]
Vector.random(10, [(10,50)]) # 10 unique numbers in [10,50]
Vector.random(30, [(10,50), 20]) # 30 unique 2D vectors
# Modes:
# 0: Unique integer vectors (default)
# 1: Non-unique integer vectors
# 2: Real-valued vectors
Vector.random(30, [(1,10), (1,10), (1,10)], 2) # 30 3D real vectors
Vector.random(30, [10], 1) # 30 numbers (may repeat)
String
Generate random text elements.
‘‘‘python
# Basic strings
String.random(5) # 5-character word
String.random((10,20), charset="abcd1234") # Variable length
String.random(10, charset="#######...") # 70% ’#’, 30% ’.’
# Structured text
String.random_sentence(5) # 5-word sentence
String.random_paragraph((3,10)) # 3-10 sentence paragraph
# Custom formatting
String.random_sentence(5, word_separators=[" "]) # Double space separator
Note: All templates support charset customization.
Sequence
Generate number sequences via recurrence.
‘‘‘python
# Explicit formula
Sequence(lambda i, f: 2*i+1) # f(i) = 2i + 1
# Recursive definition
Sequence(lambda i, f: f(i-1)+1, [0,1]) # f(i)=f(i-1)+1 with f(0)=0, f(1)=1
Sequence(lambda i, f: f(i-1)+1, {100:101, 102:103}) # Sparse base cases
# Usage
seq = Sequence(lambda i, f: f(i-1)+2, [0,2,4])
seq.get(3) # Returns 6
seq.get(4,6) # Returns [8,10,12]
Important: Recursive definitions require base cases.
Utilities
Conversion:
‘‘‘python
ati([0, 5, 100, 1E3, 1E5]) # Converts scientific notation to integers
Random Numbers:
‘‘‘python
randint(1,5) # Integer in [1,5]
uniform(1,5) # Float in [1,5]
choice([1,2,3]) # Random selection
random() # Float in [0,1)
Constants:
‘‘‘python
PI # 3.1415926...
E # 2.7182818...
ALPHABET_SMALL # "abcdefghijklmnopqrstuvwxyz"
ALPHABET_CAPITAL # "ABCDEFGHIJKLMNOPQRSTUVWXYZ"
ALPHABET # Combined letters
NUMBERS # "0123456789"

### v2

@@ 6 Discussion
Longer Reasoning Reflects Higher Task Difficulty. Table 7 shows that the pass rate decreases sharply as reasoning token length increases, exhibiting a clear downward trend. This finding runs counter to the intuitive expectation that greater test-time token usage reflects deeper reasoning and should therefore yield higher accuracy. Instead, we observe a significant chained relationship among problem difficulty, reasoning length, and pass rate: problem difficulty is positively correlated with reasoning length, while reasoning length is strongly negatively correlated with pass rate. This mediation pattern can be summarized as higher difficulty $\rightarrow$ longer reasoning length $\rightarrow$ lower pass rate.
@@
Synthetic Data Improves Test-time Scaling Efficiency. We compare the pass@k performance of Qwen2.5-Coder-7B-Instruct, Qwen3-8B, X-Coder-7B-SFT, and X-Coder-7B in Figure 6. X-Coder-7B outperforms its foundation model by 51.3 points in pass@16, and matches Qwen3-8B with 8× fewer rollouts. Moreover, X-Coder shows a larger gap between pass@1 and pass@16 compared to Qwen3-8B (19.2 vs. 13.8), indicating greater diversity in the reasoning patterns it can explore. Although RL models typically begin with higher initial performance than the SFT model, we observe that the gap does not expand within 16 rollouts, suggesting that RL improves pass@1 but might not escape its starting point largely (Wu et al., 2025).
Cognitive Behaviors Distilled from SFT, Strategic Exploitation Emerges in RL. Post-SFT models exhibit emergent cognitive behaviors like planning and verification (Appendix H.1), suggesting these capabilities are distilled directly from the teacher. In later RL stages, alongside performance gains, we also observe occasional instances of strategic exploitation, where the model games edge cases for partial rewards (Appendix H.3). We also observe persistent inefficiencies, including context-induced premature termination and cross-lingual hallucinations (e.g., translating memorized C++ solutions), as detailed in Appendix H.2.
@@ 7 Conclusion
In this paper, we explore a fully synthetic approach to competitive programming, demonstrating that rigorously verified synthetic tasks, solutions, and tests can train large reasoning models to achieve significant performance gains. Our results suggest that such synthesis offers a viable alternative to the scarcity of real-world data. Furthermore, we distill critical insights into code-centric SFT-then-RL training, shedding light on synthetic data scaling, the necessity of domain-specific evolution, and the role of code-centric reinforcement learning. We plan to release related resources.
@@
  • Section A Training and Evaluation: Reward functions, training costs, and distributed infrastructure.
  • Section B Novel Task Synthesis: Feature extraction, task generation process, difficulty and diversity analysis.
  • Section C Solution Generation and Quality Assurance: Solution validation procedures and SFT dataset statistics.
  • Section D Test Case Generation: Comparison of prompting-based vs. tool-based test generation methods.
  • Section E Dual-verification: Algorithm description, weighting criteria, and error rate analysis.
  • Section F Generality: Cross-model and cross-benchmark evaluation results.
  • Section G Data Leakage Analysis: Contamination detection and prevention measures.
  • Section H Case Study: Success and failure case analysis, reward hacking behaviors.
@@ Appendix B Novel Task Synthesis
First, rather than relying on broad feature definitions, we explicitly extract and evolve competition-related features from 10,000 question–solution pairs in TACO (Li et al., 2023) using GPT-4o-0513 (Appendix B.1). Second, we adopt a two-stage process: selecting mutually consistent features and then formulating challenging, hint-free tasks (Appendix B.2). Third, we extend the synthesis method to support multi-style generation, covering CodeForces-style tasks (rich narratives with standard I/O), LeetCode-style tasks (starter code with fixed signatures), and AtCoder-style tasks (concise specifications), thereby enhancing task diversity. In Appendix B.3, we further estimate the difficulty of synthesized problems using a trained discriminator.
@@ B.1 Feature Extraction and Evolution
While EpiCoder extracts general-purpose features from raw corpus, we explicitly extract and evolve competitive programming-related feature. Specifically, we design multiple aspect of features that highly relates to competitive programming, such as data structure, algorithm, mathematical, etc.
@@
You are provided with a feature tree represented as a nested JSON structure. Each node in this tree represents a feature or a sub-feature of competitive algorithm programming, with the leaves being the most specific features. Your task is to expand this feature tree both in depth and breadth. Depth expansion means adding more specific sub-features to existing leaves. Breadth expansion means adding more sibling features at the current levels.
@@ B.1.1 Statistics for Feature Extraction and Evolution
We present detailed statistics on feature evolution and data filtering to demonstrate how the process expands feature diversity and yields a high-quality 240k dataset. The statistics of feature extracted and evolved as follows.
@@
The evolution strategy greatly enhances diversity of features, providing support for generating diverse tasks.
@@ B.2 Stylized Task Generation for Competitive Programming
You are a professional competitive programming problem setter. Your task consists of three parts:
@@ B.2.1 Compatibale Feature Selection
We present a case to examine how model selects compatible features and combine them.
@@
LLM pairs each feature with potentially usage to obtain feature tree with role annotation. For example, LLM will anonote feature “rolling hash” as “Compute hash values for sliding windows in constant time”. These annotations help LLM to aggregate these features based on their potentially usage. For above given feature tree, the feature tree with potential usage looks like:
@@
LLM then selects a compatible and consistent subtree that can formulate a self-contained competitive programming problem. For example, LLM keeps the features that can be aggregated into selected feature tree, and concluding with an integration strategy, which displays how to combine these features into a unified problem.
@@ B.2.2 From Feature to Stylized Task
During task generation, LLM receives selected features tree and its integration strategy to formulate stylized task based on prompt recieved. In this instance, our generated Codeforces problem is shown in Figure 12.
@@
The rationale for above two-stage approach is that a single-step approach is less effective. When performing both steps simultaneously, LLMs tend to oversimplify complex instructions into trivial cases, reducing both diversity and difficulty of the generated task.
To empirically validate this, we generated 32k tasks using the one-step method (feature-tree $\rightarrow$ task) and using proposed “two-stage” method (feature-tree $\rightarrow$ sub-tree $\rightarrow$ task). The SFT results on LiveCodeBench v5 are as Table 9. The 5.3 gain shows that explicit sub-tree selection and integration is significantly helpful for producing high-quality, challenging tasks and justifies the strategy’s modular design.
Figure 13: Task style comparison.
B.2.3 Ablation on Task Style
We evaluate the effect of task styles (AtCoder, Codeforces, and LeetCode) by synthesizing three corpora of 32k tasks each (8k unique problems with 4 solutions per problem) from identical input features. For each corpus, solutions are generated with DeepSeek-R1-0528 and used to fine-tune the Qwen2.5-Coder-7B-Instruct. Results are shown in Figure 13. Although AtCoder-style tasks yield slightly higher scores, we adopt Codeforces-style as the predominant format in our demonstration dataset (Codeforces : AtCoder : LeetCode = 70 : 15 : 15), reflecting its prominence as the mainstream competitive-programming platform.
@@ B.3 Task Difficulty Estimates
Judging the difficulty of a synthetic task is challenging. To better capture the difficulty distribution of tasks generated by X-Coder, we adopt a classifier-based approach. Specifically, we add a special classification token to Qwen2.5-Coder-14B-Instruct and fine-tune it to predict the Codeforces rating of 6,246 tasks from the CodeContests dataset with annotated ratings, reserving 5% as a validation set. The fine-tuned model achieves 84% classification accuracy on the validation set. We then use this model to estimate the difficulty of 1,000 tasks generated by the proposed method, obtaining a holistic distribution as shown in Table 10.
Table 10: Difficulty distribution of Codeforces-style ratings. “Original” denotes the annotated distribution from CodeContests, and “Test” denotes 1,000 tasks generated by the proposed method.
@@ D.1 Prompting-based Test Generation
You are a professional test case generation expert, skilled at designing comprehensive test cases for programming problems. Please generate 15 different test cases for the following programming problem, including edge cases, small-scale, medium-scale, and large-scale test data.
@@ 5. Ensure test cases can thoroughly validate the correctness of solutions
D.2 Comparison of Test Generation Methods
We compare prompting-based and tool-based test generation using tasks from CodeContests (Li et al., 2022). We leverage the corresponding golden solutions to evaluate the accuracy and complexity of the tests produced by the two approaches. The results in Table 12 show that the tool-based approach outperforms the prompting-based method across multiple dimensions. Qualitatively, it is more versatile, capable of systematically generating random, scalable, boundary, and stress tests, which are essential for robust code evaluation but not supported by prompting-based methods.
Prompting-based low 13.6 5 15 82.0% 77.4%
Tool-based high 18.3 5 27 78.8% 87.9%
D.3 Tool-based Test Generation
@@ 8. The program should use argparse to provide configurable random seed control:
[The complete API documentation of the CYaRon library is provided here, covering Input/Output (IO), Graph Generation, Polygon, Vector, String, Sequence, and Utilities. We omit the full text for brevity as it follows the standard documentation of the library.]